Many years ago, in the public grid, there was an alternating voltage of the effective value 127
1.On an old lamp from that time is 127 V; 40 W. Today is the mains voltage
effective value 230
2. What would happen if this lamp were used today?
1
Expert's answer
2020-05-12T10:23:06-0400

As per the question,

The initial effective value of the voltage =127V and 40W

The resistance of the lamp "(R)=\dfrac{V^2}{P}=\dfrac{127^2}{40}"

When the effective voltage is changed to 230V, then the power of the lamp

"P=\dfrac{V^2}{R}"

Now substituting the values,

"P=\dfrac{230^2\times 40}{127^2}=131.2W"